The polynomial (my^2-4y+3)-(5y^2-2y-p) simplifies to 4y^2-2y-4 how would I find the values of m and p?
Answer by fractalier(6550)

For all of these problems, you have to look at the particular term...like I have shown you...
For (my^2-4y+3)-(5y^2-2y-p) to be equal to 4y^2-2y-4,
look at the y^2 terms...
m - 5 = 4
m = 9
and look at the constant terms...
3-(-p) = -4
3 + p = -4
p = -7